Four less than six times a number is equal to 40 more than one-half of the number.Find the number

Answer by mananth(16946)

let number be x
4 less than 6 times = 6x-4
40 more than half = 40+(x/2)
6x-4 = 40 + (x/2)
6x-(x/2) = 40+4
11x/2 = 44
x= 2*44/11
x=8
The number is 8